OpenStudy (anonymous):

WHAT IS THE AREA OF A CIRCLE WITH THE CIRCUMFERENCE OF 31.4CM?

OpenStudy (pottersheep):

Think about it. Whats the formula for circumference?

OpenStudy (anonymous):

\[circumference = 2 \pi r\]find r then \[area = \pi r^2\]

OpenStudy (anonymous):

for finding area we need value of radius .... circumference = 2πr Area = πr^2

OpenStudy (pottersheep):

31.4CM =2πr r = 4.99

OpenStudy (pottersheep):

Get it? Cuz 31.4CM divided by 2 divided by π gives you r

OpenStudy (pottersheep):

Round that number to 5.

OpenStudy (pottersheep):

Now plug the new radius into the area formula

OpenStudy (pottersheep):

πr2

OpenStudy (pottersheep):

π5^2

OpenStudy (pottersheep):

= 78.5cm3

OpenStudy (anonymous):

pottersheep has just established that:\[\text{area of a circle}=\frac{c^2}{4 \pi }\]